# Enhancing Peer Review in Astronomy: A Machine Learning and Optimization Approach to Reviewer Assignments for ALMA

### v1

@@ 6.1 Limitations
Our experience found that while the machine learning algorithms were quite successful in identifying similar proposals, a small fraction of assignments were matching proposals triggered on some individual words that are relatively rare, but can appear across different scientific categories (e.g., polarization, maser, corona). Some of these mismatches could likely be reduced with improved machine learning algorithms (e.g., using transformer-based language models and/or fine tuning to ALMA-relevant corpora), but the possibility of mismatches will always remain. In the current deployment, by using the keywords of the proposals and reviewer expertise, inappropriate assignments were flagged in advance and increased the robustness of the overall process.
6.2 Open problems
Finally, we outline several open problems in this exciting research domain. First, in assigning reviewers to the proposals, there is a tension between expertise and topology of assigned reviewers, as discussed in Section 5.4. While previous works have explored these factors independently, a promising direction is to study them jointly and establish the optimal tradeoff.
Second, unlike many other review systems that rely on ratings, ALMA requires each reviewer to rank the proposals they evaluate. Determining the most effective evaluation method is a key area of interest, closely tied to the reviewer assignment process. Rankings present challenges for cross-reviewer comparisons due to their independence, whereas ratings are subject to calibration issues such as varying strictness or leniency across reviewers (Ge et al., 2013; Roos et al., 2012; Wang & Shah, 2019). Additionally, research is ongoing into methods for eliciting both rankings and ratings and effectively integrating them (Pearce & Erosheva, 2022; Liu et al., 2022).
Third, systems like ALMA accumulate multiple cycles of reviews, presenting an opportunity to develop methods that leverage these data to continually improve reviewer assignments.
Fourth, competitive reviewing scenarios may incentivize strategic or dishonest behavior. One example is the formation of “collusion ring”, where reviewers support a paper in exchange for reciprocal favors (Littman, 2021; Vijaykumar, 2020). Colluding reviewers may manipulate the assignment system to be assigned their target paper. Some venues mitigate this risk by introducing randomization into the assignment process (Jecmen et al., 2020), limiting the likelihood of such matches. While colluders may still achieve high similarity scores, randomization reduces the probability of assignment. Another issue arises when reviewers deliberately downrate competing papers to increase the chances of their own being accepted (Balietti et al., 2016). Research in this area focuses on developing optimization methods that prevent reviewers from influencing the outcomes of their own papers (Alon et al., 2011; Dhull et al., 2022) and detecting large-scale manipulative behavior (Stelmakh et al., 2021b). Improving detection and mitigation of such behavior remains an open challenge.
As ALMA continues to refine its assignment process in future cycles, the successful deployment of AI-based methods underscores the potential for broader adoption of these technologies across the scientific community, offering a powerful tool to address both current and future challenges in peer review.

### v2

@@ 4 Summary and conclusions
Our experience also suggests that the implementation of machine learning could be improved. In a small fraction of cases, proposals were matched to reviewers based on words that are relatively rare, but can appear across different scientific categories (see Section 2.2). For example, a proposal to observe polarization in the Sun will have a finite similarity to a proposal to observe dust polarization in high-redshift galaxies since both proposals use the term “polarization.” In the current deployment, we use the keywords provided by the PIs and the reviewers to mitigate such assignments and increase the robustness of the overall process. Some of these mismatches could likely be reduced with improved machine learning algorithms (e.g., using transformer-based language models and/or fine-tuning to ALMA-relevant corpora). However, it is likely that some level of mismatch will persist (Stelmakh et al., 2023), as inherent limitations and uncertainties are a natural part of such automated processes.
The success of these methods in ALMA’s peer review process highlights the broader potential of machine learning and optimization for addressing challenges in scientific peer review, paving the way for its wider adoption across the research community.
@@ Acknowledgements
Appendix A Identifying A Reviewer’s Expertise
Section 2.2 presented the general approach to identify the expertise of each reviewer. Although this basic method works well for most reviewers, several practical issues can arise. Some reviewers are involved in proposals across distinct topics, either as PIs or coIs. Combining all their proposals into a single expertise vector may dilute their specific areas of expertise, resulting in decreased similarity when comparing against proposals focused on only one topic. Additionally, some reviewers claim expertise in a broad range of topics and submit proposals across multiple categories. Averaging their expertise based on all submitted proposals may lead to a dominance of the category that contains the most proposals. Moreover, some reviewers assert expertise in categories that differ from those of their submitted proposals, indicating a preference for reviewing in certain topics. Assuming the reviewers’ declared expertise at face value and including all proposals blindly could result in assignments outside their actual areas of expertise. Finally, there can be false positives arising from the LDA topic modeling, where a reviewer’s expertise may appear similar to a proposal topic, but upon closer inspection, the proposal may not be relevant to the reviewer’s expertise. Such false matches can occur due to common words or phrases shared across multiple categories, such as “masers” and “polarization.” This suggests that the topics identified in the LDA model may require further refinement.
To address these challenges, we utilize the keywords associated with a reviewer’s expertise to select the proposals that will be used to infer their expertise. The overall goal is to evaluate a reviewer’s expertise based on proposals that are as closely aligned as possible with the topic of the parent proposal, while also respecting the reviewer’s self-declared areas of expertise. Each ALMA proposal is submitted to a specific scientific category, defined as the parent category. We have predefined similar categories that significantly overlap with the parent category in terms of scientific content (see Figure 2). For instance, Categories 1 and 2 are considered similar because they both pertain to galaxies, while Categories 3 and 4 are similar due to their focus on the formation of stars and circumstellar disks. In contrast, Categories 1 and 2 are distinct from Categories 3, 4, and 5, and Categories 3 and 4 are distinct from Categories 1, 2, and 5. Category 5 is viewed as distinct from all other categories due to its diverse set of keywords.
With these definitions in place, a reviewer’s expertise is constructed based on whether they have provided their expertise keywords and whether they are the PI or coI on proposals. Table 1 outlines the sequence of steps used to infer the reviewer’s expertise. In the first step, if a reviewer has declared expertise on any keywords in a parent category C and has served as the PI on proposals in that category during any cycle, the expertise vector is constructed from the proposals in Category C where they are the PI. If the conditions of Step 1 are not met, Step 2 checks whether the reviewer has expertise in a similar category and has served as the PI on proposals in that category. If those conditions are also not met, Step 3 checks if the reviewer has expertise in a distinct proposal category and has served as the PI on proposals in those distinct categories. If no suitable proposals are found in these steps, the process is repeated by checking proposals in which the reviewer is a coI (Steps 4–6).
If a reviewer has not provided their expertise, Steps 7–9 examine categories in which they have submitted proposals as the PI, followed by Step 10, which considers categories where they have served as a coI in the parent category. Finally, if a reviewer has declared their expertise but has neither been a PI nor a coI on any proposals within those expertise categories, a constant similarity value is assigned for proposals in their declared expertise category.
An additional subtlety to this procedure is that there are some scientific categories (e.g., Category 5) and subtopics within categories (e.g., Solar System within Category 4) where there are a limited number of proposals and therefore a limited number of available reviewers. The above procedure could reduce the chance reviewers are assigned such proposals if they submitted proposals in other topics. If a reviewer declared they had expertise in these less popular keywords and has been PI or coI on such proposals, separate expertise vectors were computed for the parent proposal category and these other topics. The appropriate expertise vector is used when computing the similarities depending on the keyword of the proposal.
The last column in Table 1 indicates the percentage of expertise vectors that were used in each step in Cycle 10. In the vast majority of cases (94.3%), the expertise vector was computed based on the same category in which the reviewer submitted a proposal (Steps 1 and 7 combined). The remaining cases mitigated situations where the reviewer has not been a PI of proposals in the parent category.
Table 1: Steps to construct the reviewer’s expertise
Step Provided Parent Similar Distinct Percentage
Expertise category category category
1 PI 91.0%
2 PI 1.7%
3 PI 0.9%
4 coI 2.4%
5 coI 0.1%
6 coI 0.2%
7 PI 3.3%
8 PI 0.1%
9 PI 0%
10 coI 0.1%
11 0.4%